# Can Masked Autoencoders Also Listen to Birds?

### v3

@@ 3.2 Fine-tuning (M2)
Domain augmentations. To bridge the inherent domain shift between training and test data in BirdSet [^3] BirdSet training data consists of focal (directed) recordings, contrasting to test data from omnidirectional soundscapes., we utilize domain-specific augmentations while fine-tuning the encoder. While Audio-MAE also employs augmentations, it does not apply strategies tailored to bioacoustics. Our augmentations, informed by results from Rauch et al. (2024), are designed to simulate common acoustic variations in bird recordings, such as diverse background noises (i.e., noise mixing), varying signal strengths (i.e., gain mixing), and the co-occurrence of multiple vocalizations (i.e., mixup). We supplement these with spectrogram-level augmentations, including frequency and time masking. Further details on the augmentations are provided in Table 9.
@@ 3.3 Frozen Representations (M3)
Prototypical probing. We propose prototypical probing as a parameter-efficient method to leverage frozen features. This involves attaching only the prototypical pooling layer $\mathbf{f}_{\phi}$ (as described in M2) to the frozen encoder ${h}_{\beta}$. Only the parameters of the prototype vectors $\{\mathbf{p}_{c,j}\}_{j=1}^{J}$ for all classes $c$ and the final class-specific linear layers $\{g_{c}\}_{c=1}^{C}$ are trained. Similar to attentive pooling, prototypical probing utilizes the full spatial feature map $\mathbf{h}_{\beta}(x)\in\mathbb{R}^{H\times W\times D}$, preserving local structural information. This might be beneficial for bird sounds, as vocalizations typically occupy small regions of the spectrogram, where global averaging may dilute this information. Additionally, prototypical probing is parameter-efficient as the additional trainable parameters only consist of the prototypes $J\cdot C\cdot D$ and the final linear layer with a total of $J\cdot C+C$ parameters. While this scales linearly with the number of classes and prototypes, its total size remains negligible compared to the encoder. For instance, with the ViT-L encoder (approx. 300M parameters), prototypical probing for the HSN task adds only about 430k parameters (Table 5). Furthermore, it is often smaller than the overhead of attentive probing, which adds approximately $2D^{2}+D$ parameters (El-Nouby et al., 2024). Prototypical probing retains the low parameter characteristic of linear probing while efficiently exploiting non-linear spatial information crucial for discriminative performance with frozen MAE features.
@@ 4 Data and Processing
Dataset |Train| Recordings |Train| Events |Test| Segments #Classes
Pretraining
@@
Dataset. Our experiments utilize BirdSet (Rauch et al., 2024), a comprehensive benchmark for multi-label bird sound classification (i.e., classifying bird species based on their vocalizations). Unlike AudioSet, where each 10-second sample captures a wide array of sounds, bird calls are typically shorter (within 5 seconds, comparable to ESC-50 (Piczak, 2015)) and are confined to narrow frequency bands. BirdSet aggregates the training data from XC (Vellinga & Planqué, 2015), encompassing approximately 520,000 unique recordings (weakly labeled at the file level) from nearly 10,000 bird species, totaling over 3 million vocalization events. For evaluation, BirdSet provides eight downstream tasks, each consisting of a dedicated training subset and a test set derived from fully annotated soundscape recordings from different geographical regions (e.g., High Sierras Nevada (HSN) or Amazon Basin (PER)). The test sets are segmented into 5-second intervals, where each interval receives multi-label annotations indicating the presence (one or multiple) or absence of birds. This structure explicitly captures challenges like domain shift between training (focal) and test (soundscape) data, as detailed in Rauch et al. (2024). Table 2 provides a detailed overview of the datasets.
@@ 5.1 Pretraining (M1)
Figure 3: Ablations for improving the base Audio-MAE. The MAP results are reported on HSN. The $+$ symbol indicates a new component, while $\uparrow$ and $\downarrow$ denote an increase and a decrease in a parameter.
Data source. Figure 3 shows that replacing the AudioSet pretraining data with BirdSet for the base model yields a modest performance gain of 2.45%${}_{\text{p}}$ in fine-tuning. Even if extensive fine-tuning on large downstream datasets can mitigate pretraining domain mismatch, domain-specific pretraining data provides a clear advantage, especially for the performance of probing techniques (see Section 5.3). However, the most decisive gains emerge from a holistic adaptation of the entire SSL pipeline to the target domain. Thus, swapping in domain-specific data is necessary but insufficient: aligning both objective and downstream training with the structure of bird vocalizations in the domain unlocks most of the benefit.
[…]
@@ 5.2 Fine-tuning (M2)
Domain augmentations. Adapting the fine-tuning process with domain-specific data augmentations is crucial. Our sequential ablation in Figure 3 shows that applying the baseline Audio-MAE to HSN without domain-adapted augmentations yields inferior results (23.71%). Introducing domain-specific augmentations (detailed in the Appendix D) during fine-tuning provides a substantial performance uplift, increasing MAP by circa 29%${}_{\text{p}}$ over the baseline implementation. This highlights the importance of domain-aware adaptations, ensuring models can deal with the challenges in the domain data (e.g., domain shift in BirdSet).
Prototypical pooling. Replacing global averaging or the cls token with prototypical pooling further enhances classification performance when fully fine-tuning the model, setting new SOTA results. For the Bird-MAE-L model on HSN, this final modification elevates the MAP score to 55.28% from 53.15% when using global average pooling. Prototypical pooling also outperforms alternative advanced pooling mechanisms such as attentive pooling (see Table 3c). This improvement contributes to more than doubling the performance of the initial Audio-MAE baseline (23.71% MAP, before any domain specifications). As shown in Table 5, the parameter overhead of this prototypical pooling operation is minimal compared to the ViT encoder.
@@ 5.3 Frozen Representations (M3)
Prototypical probing. We ablate the quality of frozen representations using various probing methods in Table 5. We use the best-performing settings from M1 and M2, including all augmentations. All experiments use $J{=}20$ prototypes. The impact of varying $J$ is shown in Appendix F. Consistent with findings in MIM (Alkin et al., 2025), our results confirm that standard probing techniques applied to features extracted via global average pooling (linear, MLP) perform poorly with frozen MAE representations, even with the domain-specific Bird-MAE. MAP scores remain notably lower than full fine-tuning. However, methods explicitly leveraging the spatial feature map achieve considerable performance gains. Attentive probing notably improves results across model sizes. Our prototypical probing further boosts performance, outperforming attentive probing across all Bird-MAE model sizes while remaining more parameter-efficient (Table 5) with approximately 20% of parameters compared to attentive probing for this dataset. Interestingly, prototypical probing performs worse than attentive probing when applied to the general-purpose Audio-MAE, highlighting that prototype-based methods might benefit from domain-specific pretraining. For the Bird-MAE-L model, prototypical probing achieves a MAP of 49.97%, a substantial gain from MLP probing (+34.75%${}_{\text{p}}$) and only 5.31%${}_{\text{p}}$ below the full fine-tuning result (55.28%). This demonstrates that by effectively utilizing the spatial information preserved in the frozen MAE feature map, prototypical probing addresses the limitations of frozen MIM representations for discriminative tasks. It offers an efficient alternative to full fine-tuning.
ViT Technique Linear MLP Attentive Prototypical
Fine-tuning: 43.11
@@
Table 4: Frozen representation ablations on HSN, evaluated with probing techniques. Linear and MLP utilize the global average, attentive and prototypical ($J\!=20$) the feature map.
Probing Parameters HSN
ViT-L parameters: 307M
Linear $C(D+1)$ $21$k
MLP $H(D+1)+C(H+1)$ $535$k
Attentive $2D^{2}+(C+1)D+C$ $2.1$M
Prototypical $C\cdot\bigl{[}J(D+1)+1\bigr{]}$ $430$k
@@ 6 Benchmark Results
Baselines and evaluation. We compare Bird-MAE against several relevant models. Our baseline is the Audio-MAE-Base [^4] Larger model checkpoints are not available from the source paper (Huang et al., 2022). pretrained on AudioSet and fine-tuned with prototypical pooling. First, we compare against other bird-specific SSL models: BirdAVES (Hagiwara, 2023) and a SimCLR (Chen et al., 2020b) implementation from Moummad et al. (2024), both pretrained on custom XC data. Second, we include results from the best-performing supervised models in bird sound classification: Google’s Perch (Hamer et al., 2023) and BirdSet’s BirdNeXt (Rauch et al., 2024), also pretrained on XC data. We report Bird-MAE results using ViT-Base, ViT-Large, and ViT-Huge backbones, incorporating all modifications from Section 3. Our domain-specific augmentation pipeline is applied during fine-tuning and probing to all models where feasible to ensure fair comparison in both tasks. We exclude spectrogram-level augmentations for the waveform-based BirdAVES model. For Perch, which is not publicly available for fine-tuning, we follow the inference protocol of masking logits for classes not present in the downstream task’s label set (Rauch et al., 2024). Hyperparameters for all models are tuned once on HSN before final evaluation across downstream tasks. More details and hyperparameters can be found in Appendix D.
@@ 6.1 Multi-Label Classification
What is the performance gain of a domain-specific MAE? Table 6 confirms results from our ablation studies: Domain specification via Bird-MAE yields substantial performance improvements across the BirdSet benchmark compared to the general-purpose Audio-MAE. While the Bird-MAE-B also offers notable gains over the available Audio-MAE baseline, the benefits become more pronounced with larger architectures. For instance, our Bird-MAE-L achieves notably higher MAP scores than the baseline, with performance gains of +15%${}_{\text{p}}$ on POW (+7%${}_{\text{p}}$ Bird-MAE-B) or +13%${}_{\text{p}}$ on PER (+6%${}_{\text{p}}$ Bird-MAE-B). Furthermore, Bird-MAE consistently and notably outperforms the other domain-specific SSL baselines AVES and SimCLR across all datasets, often by margins exceeding 15-20%${}_{\text{p}}$ MAP on average.
How does the model compare to supervised models? We compare Bird-MAE against the current best-performing supervised models (Perch and BirdNeXt). Our fine-tuned Bird-MAE models consistently achieve new SOTA results across the BirdSet benchmark. Specifically, Bird-MAE-L outperforms the BirdNeXt and Perch baselines on all eight datasets, often by considerable margins. For example, on SSW, Bird-MAE-L achieves 40.82% MAP compared to Perch’s 28.11% MAP (+12.7%${}_{\text{p}}$), and on PER, it achieves 34.64% MAP versus Perch’s 18.77% MAP (+15.9%${}_{\text{p}}$). These results demonstrate the effectiveness of domain-specific SSL combined with modern transformer architectures and higher parameter counts compared to prior supervised CNN-based approaches in bird sound classification.
@@
Supervised
Perch EffNet-B1 Xeno-Canto 41.84 41.12 18.77 39.12 27.81 63.62 28.11 29.45
BirdNeXt ConvNeXt XCL BirdSet 45.72 37.97 20.98 31.84 22.69 66.56 33.23 26.13
Self-Supervised
@@
Can we freeze the representations? While audio SSL currently relies on fine-tuning (Chen et al., 2023; 2024a), efficient deployment is highly desirable for edge applications in bioacoustics (Höchst et al., 2022). We evaluate the performance of frozen representations using linear versus our proposed prototypical probing in Table 7. Linear probing performs poorly for Audio-MAE and Bird-MAE across backbone sizes, confirming the difficulty of using frozen MIM representations directly. However, prototypical probing drastically improves performance by leveraging the spatial feature map: it closes the gap to fine-tuning to approximately 3%${}_{\text{p}}$ MAP on average across downstream tasks. Bird-MAE with prototypical probing also achieves substantial gains over Audio-MAE with prototypical probing (e.g., +27.3%${}_{\text{p}}$ base performance on NBP). Additionally, it notably outperforms other SSL models (AVES, SimCLR) using either probing method and also surpasses the fully supervised Perch model (from Table 6) on nearly all datasets using only frozen representations. This challenges the notion that MAE features are unsuitable for probing and demonstrates prototypical probing as a highly effective and efficient alternative to fine-tuning in bird sound classification. While prototypical probing enhances performance for the masking-based BirdAVES model, it seems to degrade performance for the contrastive SimCLR model, suggesting probing effectiveness interacts with the SSL pretraining objective.
@@ 6.2 Few-Shot Multi-Label Probing
Settings. In this section, we evaluate the few-shot learning capabilities of Bird-MAE, introducing a few-shot multi-label classification benchmark in BirdSet. This setup maintains the standard test sets and domain shifts but restricts the training data to $k\in\{1,5,10\}$ event instances per class for each downstream task’s training subset. We only report our best-performing Bird-MAE-L model with an average of over three repetitions and three randomly sampled subsets per shot. We focus on the performance of SSL models’ frozen representations in low data regimes. Detailed results are available in Appendix E.2, and the few-shot sampling strategy is described in more detail in Appendix A.
@@
7 Conclusion and Future Work

### v4

@@ 3.2 Fine-tuning (M2)
Domain augmentations. To bridge the inherent domain shift between training and test data in BirdSet [^3] BirdSet training data consists of focal (directed) recordings, contrasting to test data from omnidirectional soundscapes., we utilize domain-specific augmentations while fine-tuning the encoder. While Audio-MAE also employs augmentations, it does not apply strategies tailored to bioacoustics. Our augmentations, informed by results from Rauch et al. (2024b), are designed to simulate common acoustic variations in bird recordings, such as diverse background noises (i.e., noise mixing), varying signal strengths (i.e., gain mixing), and the co-occurrence of multiple vocalizations (i.e., mixup). We supplement these with spectrogram-level augmentations, including frequency and time masking. Further details on the augmentations are provided in Table 9.
@@ 3.3 Frozen Representations (M3)
Prototypical probing. We propose prototypical probing as a parameter-efficient method to leverage frozen features. This involves adapting the prototypical pooling layer $\mathbf{f}_{\phi}$ (as described in M2) to the frozen encoder ${h}_{\beta}$ as a lightweight probing head. It is a parametric method: The parameters of the prototype vectors $\{\mathbf{p}_{c,j}\}_{j=1}^{J}$ for all classes $c$ and the final class-specific linear layers $\{g_{c}\}_{c=1}^{C}$ are trained. Similar to attentive pooling, prototypical probing utilizes the full spatial feature map $\mathbf{h}_{\beta}(x)\in\mathbb{R}^{H\times W\times D}$, preserving local structural information. This might be beneficial for bird sounds, as vocalizations typically occupy small regions of the spectrogram, where global averaging may dilute this information. Additionally, prototypical probing is parameter-efficient as the additional trainable parameters only consist of the prototypes $J\cdot C\cdot D$ and the final linear layer with a total of $J\cdot C+C$ parameters. While this scales linearly with the number of classes and prototypes, its total size remains negligible compared to the encoder. For instance, with the ViT-L encoder (approx. 300M parameters), prototypical probing for the HSN task adds only about 430k parameters (Table 5.3). Furthermore, it is often smaller than the overhead of attentive probing, which adds approximately $2D^{2}+D$ parameters (El-Nouby et al., 2024). Prototypical probing retains the low parameter characteristic of linear probing while efficiently exploiting non-linear spatial information crucial for discriminative performance with frozen MAE features.
@@ 4 Data and Processing
[…]
@@
Dataset. Our experiments utilize BirdSet (Rauch et al., 2024b), a comprehensive benchmark for multi-label bird sound classification (i.e., classifying bird species based on their vocalizations). Unlike AudioSet, where each 10-second sample captures a wide array of sounds, bird calls are typically shorter (within 5 seconds, comparable to ESC-50 (Piczak, 2015)) and are confined to narrow frequency bands. BirdSet aggregates the training data from XC (Vellinga & Planqué, 2015), encompassing approximately 520,000 unique recordings (weakly labeled at the file level) from nearly 10,000 bird species, totaling over 3 million vocalization events. For evaluation, BirdSet provides eight downstream tasks, each consisting of a dedicated training subset and a test set derived from fully annotated soundscape recordings from different geographical regions (e.g., High Sierras Nevada (HSN) or Amazon Basin (PER)). The test sets are segmented into 5-second intervals, where each interval receives multi-label annotations indicating the presence (one or multiple) or absence of birds. This structure explicitly captures challenges like domain shift between training (focal) and test (soundscape) data, as detailed in Rauch et al. (2024b). Table 2 provides a detailed overview of the datasets.
@@ 5.1 Pretraining (M1)
Figure 3: Ablations for improving the base Audio-MAE. The MAP results are reported on HSN. The ++ symbol indicates a new component, while $\uparrow$ and $\downarrow$ denote an increase and a decrease in a parameter.
Data source. Figure 3 shows that replacing the AudioSet pretraining data with BirdSet for the base model yields a modest performance gain of 2.45 pp in fine-tuning. Even if extensive fine-tuning on large downstream datasets can mitigate pretraining domain mismatch, domain-specific pretraining data provides a clear advantage, especially for the performance of probing techniques (see Section 5.3). However, the most decisive gains emerge from a holistic adaptation of the entire SSL pipeline to the target domain. Thus, swapping in domain-specific data is necessary but insufficient: aligning both objective and downstream training with the structure of bird vocalizations in the domain unlocks most of the benefit.
Data processing. The quality and size of the pretraining dataset are crucial for SSL. To quantify the impact of dataset size, we begin with the full $\text{{XCL-3.4M}}_{\text{{R}}}$ training dataset with all available sound events and progressively reduce it to 50%, 25%, and 12.5% by random sampling. As shown in Table 3a, pretraining performance generally scales with dataset size when using randomly sampled subsets, although gains diminish beyond 1.7 million samples in our use case. Noticeably, applying our data curation strategy (balancing classes, reducing redundancy via metadata) to create the curated 1.7 million sample dataset (XCL-1.7M) results in better performance compared to using an uncurated dataset of the same size or even the full, uncurated 3.4 million sample dataset. This highlights the benefit of data curation, outweighing data volume in this setting. More details of the curation process are available in the Appendix A.
@@ 5.2 Fine-tuning (M2)
Domain augmentations. Adapting the fine-tuning process with domain-specific data augmentations is crucial. Our sequential ablation in Figure 3 shows that applying the baseline Audio-MAE to HSN without domain-adapted augmentations yields inferior results (23.71%). Introducing domain-specific augmentations (detailed in the Appendix D) during fine-tuning provides a substantial performance uplift, increasing mAP by circa 29 pp over the baseline implementation. This highlights the importance of domain-aware adaptations, ensuring models can deal with the challenges in the domain data (e.g., domain shift in BirdSet).
Prototypical pooling. Replacing global averaging or the cls token with prototypical pooling further enhances classification performance when fully fine-tuning the model, setting new SOTA results. For the Bird-MAE-L model on HSN, this final modification elevates the mAP score to 55.28% from 53.15% when using global average pooling. Prototypical pooling also outperforms alternative advanced pooling mechanisms such as attentive pooling (see Table 3c). This improvement contributes to more than doubling the performance of the initial Audio-MAE baseline (23.71% mAP, before any domain specifications). As shown in Section 5.3, the parameter overhead of this prototypical pooling operation is minimal compared to the ViT encoder.
@@ 5.3 Frozen Representations (M3)
Prototypical probing. We ablate the quality of frozen representations using various probing methods in Section 5.3. We use the best-performing settings from M1 and M2, including all augmentations. All experiments use J=20J{=}20 prototypes. The impact of varying $J$ is shown in Appendix F. Consistent with findings in MIM (Alkin et al., 2025), our results confirm that standard probing techniques applied to features extracted via global average pooling (linear, MLP) perform poorly with frozen MAE representations, even with the domain-specific Bird-MAE. mAP scores remain notably lower than full fine-tuning. However, methods explicitly leveraging the spatial feature map achieve considerable performance gains. Attentive probing notably improves results across model sizes. Our prototypical probing further boosts performance, outperforming attentive probing across all Bird-MAE model sizes while remaining more parameter-efficient (Section 5.3) with approximately 20% of parameters compared to attentive probing for this dataset. Interestingly, prototypical probing performs worse than attentive probing when applied to the general-purpose Audio-MAE, highlighting that prototype-based methods might benefit from domain-specific pretraining. For the Bird-MAE-L model, prototypical probing achieves a mAP of 49.97%, a substantial gain from MLP probing (+34.75 pp) and only 5.31 pp below the full fine-tuning result (55.28%). This demonstrates that by effectively utilizing the spatial information preserved in the frozen MAE feature map, prototypical probing addresses the limitations of frozen MIM representations for discriminative tasks. It offers an efficient alternative to full fine-tuning.
[…]
@@
Table 4: Frozen representation ablations on HSN, evaluated with probing techniques. Linear and MLP utilize the global average, attentive and prototypical (J=20J\!=20) the feature map.
[…]
Linear $C(D+1)$ 2121k
MLP $H(D+1)+C(H+1)$ 535535k
Attentive $2D^{2}+(C+1)D+C$ 2.12.1M
Prototypical $C\cdot\bigl{[}J(D+1)+1\bigr{]}$ 430430k
@@ 6 Benchmark Results
Baselines and evaluation. We compare Bird-MAE against several relevant models. Our baseline is the Audio-MAE-Base [^4] Larger model checkpoints are not available from the source paper (Huang et al., 2022). pretrained on AudioSet and fine-tuned with prototypical pooling. First, we compare against other bird-specific SSL models: BirdAVES (Hagiwara, 2023) and a SimCLR (Chen et al., 2020b) implementation from Moummad et al. (2024), both pretrained on custom XC data. Second, we include results from the best-performing supervised models in bird sound classification: Google’s Perch (Hamer et al., 2023) and BirdSet’s BirdNeXt (Rauch et al., 2024b), also pretrained on XC data. We report Bird-MAE results using ViT-Base, ViT-Large, and ViT-Huge backbones, incorporating all modifications from Section 3. Our domain-specific augmentation pipeline is applied during fine-tuning and probing to all models where feasible to ensure fair comparison in both tasks. We exclude spectrogram-level augmentations for the waveform-based BirdAVES model. For Perch [^5] Perch is not publicly available for fine-tuning. and BirdNeXt, we mask logits for classes not present in the downstream task’s label set (Rauch et al., 2024b). Hyperparameters for all models are tuned once on HSN before final evaluation across downstream tasks. More details and hyperparameters can be found in Appendix D.
@@ 6.1 Multi-Label Classification
[…]
@@
What is the performance gain of a domain-specific MAE? Table 6 confirms results from our ablation studies: Domain specification via Bird-MAE yields substantial performance improvements across the BirdSet benchmark compared to the general-purpose Audio-MAE. While the Bird-MAE-B also offers notable gains over the available Audio-MAE baseline, the benefits become more pronounced with larger architectures. For instance, our Bird-MAE-L achieves notably higher mAP scores than the baseline, with performance gains of +15 pp on POW (+7 pp Bird-MAE-B) or +13 pp on PER (+6 pp Bird-MAE-B). Furthermore, Bird-MAE consistently and notably outperforms the other domain-specific SSL baselines AVES and SimCLR across all datasets, often by margins exceeding 15-20 pp mAP on average.
How does the model compare to supervised models? We compare Bird-MAE against the current best-performing supervised models (Perch and BirdNeXt). Our fine-tuned Bird-MAE models consistently achieve new SOTA results across the BirdSet benchmark. Specifically, Bird-MAE-L outperforms the BirdNeXt and Perch baselines on all eight datasets, often by considerable margins. For example, on SSW, Bird-MAE-L achieves 40.82% mAP compared to Perch’s 28.11% mAP (+12.7 pp), and on PER, it achieves 34.64% mAP versus Perch’s 18.23% mAP (+16.4 pp). These results demonstrate the effectiveness of domain-specific SSL combined with modern transformer architectures and higher parameter counts compared to prior supervised CNN-based approaches in bird sound classification.
Can we freeze the representations? While audio SSL currently relies on fine-tuning (Chen et al., 2023; 2024a), efficient deployment is highly desirable for edge applications in bioacoustics (Höchst et al., 2022). We evaluate the performance of frozen representations using linear versus our proposed prototypical probing in Table 7. Linear probing performs poorly for Audio-MAE and Bird-MAE across backbone sizes, confirming the difficulty of using frozen MIM representations directly. However, prototypical probing drastically improves performance by leveraging the spatial feature map: it closes the gap to fine-tuning to approximately 3 pp mAP on average across downstream tasks. Bird-MAE with prototypical probing also achieves substantial gains over Audio-MAE with prototypical probing (e.g., +27.3 pp base performance on NBP). Additionally, it notably outperforms other SSL models (AVES, SimCLR) using either probing method and also surpasses the fully supervised Perch model (from Table 6) on nearly all datasets using only frozen representations. This challenges the notion that MAE features are unsuitable for probing and demonstrates prototypical probing as a highly effective and efficient alternative to fine-tuning in bird sound classification. While prototypical probing enhances performance for the masking-based BirdAVES model, it seems to degrade performance for the contrastive SimCLR model, suggesting probing effectiveness interacts with the SSL pretraining objective.
@@ 6.2 Few-Shot Multi-Label Probing
Settings. In this section, we evaluate the few-shot learning capabilities of Bird-MAE, introducing a few-shot multi-label classification benchmark in BirdSet to test the frozen representations in a low data regime. This setup maintains the standard test sets and domain shifts but restricts the training data to $k\in\{1,5,10\}$ event instances per class for each downstream task’s training subset. We only report our best-performing Bird-MAE-L model with an average of three repetitions and three randomly sampled subsets per shot. Detailed results are available in Appendix E.2 and the few-shot sampling strategy is described in Appendix A.
@@
7 Discussion and Limitations
Several research directions emerge from our findings, framed by the scope and limitations of this study. In the following, we discuss these limitations and show important directions for future work.
Upfront cost of domain-specific pretraining. We advocate for a domain-specific foundation model in complex fields like bioacoustics. This approach requires a notable one-time investment for pretraining, which we frame as a necessary trade-off for achieving SOTA performance in challenging fine-grained domains where general-purpose models currently fall short (Hamer et al., 2023; Turian et al., 2022). This initial cost is balanced by a reusable asset for the research community that enables efficient downstream adaptation (Ghani et al., 2023): Various sub-tasks can be tackled by simply training a lightweight probe, avoiding the need for repeated, costly fine-tuning. For instance, Bird-MAE could be deployed for other bioacoustic tasks like fine-grained call-type classification (Kahl et al., 2021) or population density estimation (Navine et al., 2024). Furthermore, our preliminary results in Table 8 suggest this benefit may even possess cross-species transferability within bioacoustics, as Bird-MAE outperforms the general-purpose model on the MeerKAT dataset. Future work could focus on making this domain-specific pretraining more computationally efficient.
[…]
Audio-MAE Linear 13.19 37.3
Audio-MAE Proto 17.97 38.6
BEATs Linear 38.3
(a) AS-20k
[…]
Audio-MAE ViT-B Linear 16.16
Audio-MAE ViT-B Proto 28.30
Bird-MAE ViT-B Linear 17.38
Bird-MAE ViT-B Proto 34.58
(b) MeerKAT
Table 8: Generalizability of prototypical layers and Bird-MAE (mAP%). Performance comparison on two non-bird datasets: general audio (AS-20k) and another fine-grained bioacoustic task (MeerKAT). We evaluate the impact of using our prototypical head versus a standard linear head for both frozen feature probing and full fine-tuning.
Recipe and results transferability. While our pretraining recipe (e.g., extended epochs, mixup) and prototypical probing are holistically adapted for bird bioacoustics, we do not evaluate their transferability to general audio tasks. To investigate the broader applicability of prototypical probing, we present preliminary experiments on two non-bird datasets: the general audio benchmark AS-20k and another fine-grained bioacoustic task, the MeerKAT mammal vocalization dataset. Details on the MeerKAT dataset are provided in Appendix G. Table 8 provides initial evidence of generalizability. On general audio (AS-20k), simply replacing the standard linear head with our prototypical pooling layer notably improves the performance of the off-the-shelf Audio-MAE for both frozen feature probing and full fine-tuning. This fine-tuning result surpasses even the more advanced BEATs (Chen et al., 2023). On the related bioacoustic MeerKAT task, we observe two key results: First, prototypical probing again substantially outperforms linear probing for both the general Audio-MAE (+12.1 pp) and our Bird-MAE (+17.2 pp). Second, Bird-MAE outperforms the general Audio-MAE with both probes, supporting findings from Ghani et al. (2023) that bioacoustic pretraining can offer benefits across related fine-grained animal vocalization tasks. These results suggest that our prototype-based methods are not limited to bird sound classification and can unlock greater performance from existing SSL backbones. A comprehensive cross-domain study and extending these methods to other fine-grained settings such as insect call or speaker dialect classification remain important next steps.